You have calculated the following holding period returns for an individual stock:

Periods               Return

Month 1                 7%

Month 2                  14%

Month 3                 -1%

What is the stock's standard deviation? Assume we don't know the probabilities about the future.

* Hint: Please calculate E(R) first and calculate standard deviation accordingly.

Please round your answer to the fourth decimal. If your answer is 1.11%, you should input "0.0111".
Answer #1
X (X - X̄) (X - X̄)²
0.07 0.00 0.00
0.14 0.07 0.01
-0.01 -0.08 0.01

mean =    ΣX/n =    0.2/3=   0.07

sample std dev =   √ [ Σ(X - X̄)²/(n-1)] =   √(0.0113/2)=   0.0751

Please let me know in case of any doubt.

Thanks in advance!


Please upvote!